A stock has an expected return of 12.60%, the risk-free rate is 3.55%, the market risk premium is 8.10%. What must be the beta of this stock ?
Answer #1

Expected return on the stock formula:

\(E R=R_{f}+\beta\left(R_{m}-R_{f}\right)\)

Where, \(\mathrm{ER}=\) Expected return 

\(\mathrm{Rf}=\) Risk-free rate 

\(\beta=\) Beta 

\(\mathrm{Rm}=\) Expected return on the market (Rm - Rf) = Market risk premium

Therefore,

\(0.1260=0.0355+\beta(0.081)\)

\(\beta(0.081)=0.1260-0.0355\)

\(\beta(0.081)=0.0905\)

\(\beta=\frac{0.0905}{0.081}\)

\(=1.11728\)

OR

\(=1.12..._{\text { (rounding up to } 2 \text { decimals) }}\)

Therefore, the beta of this stock is 1.12 .


answered by: Marie